if four congruent circles are insicribed in a sqaure qith the side length of 20mm what is the peremiter of the shaded portion of the square?: if four congruent circles are insicribed in a sqaure qith the side length of 20mm what is the peremiter of the shaded portion of the square?
Answer by mangopeeler07(445)

if the side of the square is 20mm, then the diameter of each circle is 20mm. That makes the radius of each circle 10mm. Now find the area of the circles (which would overlap in a square, so I guess I should say the circle).

area of circle=pi(r^2)
area of circle=pi(10^2)
area of circle=100pi

If the circle is shaded, then that is the area of the shaded part of the square.

Area of shaded region= 100(pi)mm^2

If everything but the circles are shaded, find the area of the square.

area of square=L^2
area of square=20^2
area of square=400

Now subtract the circle from that.

Area of shaded region=400-100(pi)mm^2